I got this wrong, and the instructor did not explain why. Can you enlighten me?
There are 385 surfers at Surf City. 2/3 were boys and 1/12 were girls. If there are 2 girls for every boy, how many boys and how many girls are there in Surf City?
I got 385(.66) = 130.90
1/12(2)(385)=247
That only equals 377 in lieu of the 385

Answer by checkley71(8403)

2/3*385=256.7
1/12*385=32.08
THIS TOTALS ONLY 289 SURFERS.
IF THE NUMBER OF GIRLS=2 TIMES THE NUMBER OF BOYS THEN:
G+B=385 & FOR EVERY BOY THERE ARE 2 GIRLS THUS B=2G. THUS:
G+2G=385
3G=385
G-385/3
G=128.333 GIRLS SOMETHING IS MISSING HERE????????????????